OpenStudy (anonymous):

simplify the sqaure root of 5/3

OpenStudy (anonymous):

\[3\sqrt{^5}\]

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

Written in radical form?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

okay then what I already said^ \[3\sqrt{^5}\]

OpenStudy (anonymous):

Do you know how I got that? @oliviamiller4

OpenStudy (anonymous):

no could you please help?

OpenStudy (anonymous):

@TaylorS11

OpenStudy (anonymous):

okay the first part of a radical expression is the denominator \[denominator \sqrt{2}\]

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

i think sothis is what the question looked like

OpenStudy (anonymous):

@TaylorS11

OpenStudy (solomonzelman):

Multiply top and bottom times cube root of 9.

OpenStudy (anonymous):

exactly what @SolomonZelman said

OpenStudy (solomonzelman):

(if 3 is in the root )

OpenStudy (anonymous):

was it written like this: \[\sqrt{3^5}\]

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \sqrt[3]{\frac{5}{3}}=\frac{\sqrt[3]{5}}{\sqrt[3]{3}}=\frac{\sqrt[3]{5}\color{blue}{\times\sqrt[3]{9}} }{\sqrt[3]{3}\color{blue}{\times\sqrt[3]{9}}} ==\frac{\sqrt[3]{45}}{3}}\)

OpenStudy (anonymous):

no it was written as a fraction @SolomonZelman

OpenStudy (anonymous):

just like this